OpenStudy (jaweria):

I need help with 11 questions of Derivatives problems in Calculus 1. Can anyone help me with that please let me know, I will really appreciate it.

OpenStudy (curry):

just post, i got you.

OpenStudy (jaweria):

ok first one is: (2x-3)^2+(4y-5)^2=10

OpenStudy (jaweria):

and we need to find dy/dx for all of them.

OpenStudy (anonymous):

you can either use implicit differentiation from the start, or you can begin by isolating the y variable.

OpenStudy (jaweria):

second question is: x^2sin(x)+y^2cos(y)=1

OpenStudy (anonymous):

likewise for the second

OpenStudy (jaweria):

yeah we have to do implicit diffentiation but I am really stuck in all of them.

OpenStudy (anonymous):

i'll get you started with (2x-3)^2+(4y-5)^2=10

OpenStudy (jaweria):

ohh thank you so much

OpenStudy (anonymous):